OpenStudy (anonymous):

Not sure where to start... find the derivative of y = sin(sinx + cosx). Can you mention rules used please?

OpenStudy (anonymous):

\(y = \sin(\sin(x) + cos(x))\), right?

OpenStudy (anonymous):

yep that's it!

OpenStudy (anonymous):

We can use chain rule here very well.. :)

OpenStudy (anonymous):

\[y' = \frac{d}{dx}\sin(\sin(x) + \cos(x)) \times \frac{d}{dx}(\sin(x) + \cos(x))\]

OpenStudy (anonymous):

Firstly, we have to take derivative of whole \(sin(w)\), then we do the derivative of \(w\), where \(w = sin(x) + cos(x)\)

OpenStudy (anonymous):

Getting?

OpenStudy (anonymous):

good so far! (ignore when it says I'm typing, my comp is effed up haha)

OpenStudy (anonymous):

The same it does with me, so don't take it granted that when you see that, then it means I am typing only.. :P

OpenStudy (anonymous):

So, what is the derivative of \(\sin(x)\), can you tell?

OpenStudy (anonymous):

cosx, right? so would the whole derivative be... (cos(cosx-sinx))(cosx-sinx) ?

OpenStudy (anonymous):

Wait.. ")

OpenStudy (anonymous):

Don't go with the speed of light, we will check it one by one, alas.!! you are not right with the whole derivative.. :( But you are right about cos(x).. :)

OpenStudy (anonymous):

This way the derivative of \(sin(w)\) will be \(cos(w)\), right?

OpenStudy (anonymous):

So, we are now having: \[y' = \cos(\sin(x) + \cos(x)) \times \frac{d}{dx}(\sin(x) + \cos(x))\]

OpenStudy (anonymous):

Getting?

OpenStudy (anonymous):

we will firstly differentiate sin only, but not its angle, for angle we have second step.. :)

OpenStudy (anonymous):

mmkay, so even though sinx's derive is cos, the values inside (w) will remain the same even if they are sin and cos?

OpenStudy (anonymous):

They are angle or you can say they are argument, see: \[\sin(\frac{\pi}{3})\] Then agnle pi/3 is just same as (sin(x)+cos(x) there..

OpenStudy (anonymous):

Remember this: first the function, then its angle.. For trigonometric case.. :)

OpenStudy (anonymous):

ahh alright, I see!

OpenStudy (anonymous):

*angle not agnle, there.. :)

OpenStudy (anonymous):

So, can we proceed further?

OpenStudy (anonymous):

yes, I'm ready!

OpenStudy (anonymous):

\[\frac{d}{dx}(\sin(x) + \cos(x)) = \frac{d}{dx} \sin(x) \times \frac{d}{dx}(x) + \frac{d}{dx}(\cos(x)) \times \frac{d}{dx}(x)\]

OpenStudy (anonymous):

This is for the second part, see first the function and then its angle.. Getting?

OpenStudy (anonymous):

hmm, okay. I think so, but we're also deriving the (x)'s? that will just be 1, right?

OpenStudy (anonymous):

Yep, that is why sin(x) derivative is cos(x): It is actually like this: \[\frac{d}{dx}(sin(x)) = cos(x) \times \frac{d}{dx}(x) = cos(x) \quad only.\]

OpenStudy (anonymous):

I am just showing it to prove that we differentiate function like sin, cos etc then we take derivative of its angle.. :)

OpenStudy (anonymous):

oh okay, I get it then!

OpenStudy (anonymous):

So, can you tell now what is our whole derivative?

OpenStudy (anonymous):

(sorry, my computer hates this website haha) y1 = (cos(sinx+cosx))(cosx-sinx)?

OpenStudy (anonymous):

Superb.. :)

OpenStudy (anonymous):

yay thank you so much for all your help!

OpenStudy (anonymous):

It will be more nice if you bring out (cos(x)-sin(x)) in the front..

OpenStudy (anonymous):

\[\color{blue}{y' = (\cos(x)-\sin(x)) \cdot \cos(\sin(x) + \cos(x))}\]

OpenStudy (anonymous):

In multiplication, you can change the order which you are writing in.. :)

OpenStudy (anonymous):

ahh I see, excellent! thanks again!! (:

OpenStudy (anonymous):

\(\text{This is it}...\) you are welcome dear.. :)